General Physics

A hollow sphere and a solid sphere having same mass and same radii are rolled down a rough inclined plane.

priya
12 Years agoGrade

1 Answer

Saurabh Koranglekar
6 Years ago
The solid sphere reaches the bottom with greater speed.
Explanation: For a hollow sphere, the linearacceleration
a= (3/5)g.sinθ= 0.60g.sinθ
For solid sphere a = (5/7)g.sinθ= 0.71g.sinθ
Since 'a' forsolidsphere is more so it will reach the bottom witha greaterspeed. Since both the spheres are released from the same position, their initial potential energymghwill be the same. So when they reach the bottom they losesameamount of P.E. and this loss of P.E. will be the gain in K.E. So both the spheres will have same K.E.at the bottom.